OpenStudy (anonymous):

Here is a list of what John weighed in pounds on six different days. Find the mean of his weight to the nearest pound. 98, 101, 96, 95, 100, 96 A. 96 pounds B. 97 pounds C. 100 pounds D. 98 pounds

OpenStudy (diyadiya):

97.6

OpenStudy (anonymous):

(98+101+96+95+100+96)/6 = 586/6= 97.777 (98 pounds

OpenStudy (anonymous):

so D?

OpenStudy (anonymous):

yes

OpenStudy (diyadiya):

yup